OpenStudy (anonymous):

If the period of a pendulum on another planet is four times its period measured on earth, what is the gravitational acceleration on the planet. Solve, and what is the equation you will use for the calculations?

OpenStudy (anonymous):

\[T_0=\sqrt{\frac{l}{g}}\]

OpenStudy (anonymous):

\[(T_0 )^2= \frac{l}{g}\] \[(4T_0 )^2= \frac{l}{g}\] \[(4T_0 )^2= \frac{l}{g}\] \[(T_0 )^2= \frac{l}{16 g}\]

OpenStudy (anonymous):

awesome, so what would it be on earth 1/4g?